Explain how to find whether 10mi/h is faster then 15ft/s.
Found 2 solutions by Fombitz, mananth:
Answer by Fombitz(32388)

Convert from one units set into another.
Either convert mph to ft/s or ft/s to mph.
To convert from X mph,
X mi/hr*(5280 ft/mi)*(1 hr/3600 sec)=Y ft/sec
or
X ft/s*(1 mi/5280 ft)*(3600 sec/1 hr)=Z mph
Work it through and find out.

Answer by mananth(16949)

10 mi /hour
= 10*5280 feet / 3600 seconds
=52,800/3600
=14.66 ft/second.
it is less